OpenStudy (anonymous):

How do I graph sin/cos functions using points for: y=3/2cos(x+pi/2)

OpenStudy (anonymous):

For x, plug in values like: -pi/2 , 0, pi/2, pi, 3pi/2 And see what the corresponding values of y turn out to be for each of those x's Plot those points and connect the dots with a smooth curve.

OpenStudy (campbell_st):

the general form is y = acos(bx + c) a is the amplitude the period of the curve is determined by 2pi/b c is the phase shift... so in your question a = 3/2 so thats the amplitude b = 1 so the period is 2pi ( it takes a distance of 2 pi to complete a full cycle) c = pi/2 so the cos curve is shifted pi/2 to the left. hope it helps

OpenStudy (anonymous):

@campbell_st I did understand it, thing is, it requires me to give exact points and was not able to just provide that answer, but thank you. @BangkokGarrett that makes complete sense, I just hope it does work once I use the program to graph it. thanks!

OpenStudy (anonymous):

Okay so I tried the first way, and was not any use since in the program I cannot use their calculator. Therefore, I need to find the points myself, how can I find the points using a  this T chart